# Extremely luminous optical afterglow of a distant and energetic gamma-ray burst GRB 230204B

### v1

@@ Abstract
Robotic telescope networks play an important role in capturing early and bright optical afterglows, providing critical insights into the energetics and emission mechanisms of GRBs. In this study, we analyze GRB 230204B, an exceptionally energetic and multi-pulsed long GRB, detected by the Fermi GBM and MAXI detectors, with an isotropic equivalent gamma-ray energy exceeding 1054 erg. Time-resolved spectral analysis reveals a transition in the prompt emission from hard (sub-photospheric dominated) spectra during early pulses to softer (synchrotron radiation dominated) spectra in later pulses, indicative of a hybrid jet composition. We report the discovery and characterization of the optical afterglow using the MASTER and BOOTES robotic telescope networks, alongside long-term radio observations extending to 335 days post-burst with the ATCA. At 1.3 ks post-burst, the optical luminosity was exceptionally high, surpassing even other bright GRBs, such as GRB 221009A (the “BOAT”). Multi-wavelength modeling, incorporating data from MASTER, BOOTES, DOT, Swift/XRT, and radio observations, was conducted using an external ISM forward-shock top-hat jet model with afterglowpy. The results reveal a narrow and highly collimated jet with a circumburst density of 28.12 cm-3, kinetic energy 4.18 erg, and a relatively low value of = 2.14 , indicating shock-compression of magnetic field in the surrounding interstellar medium. We constrained a low radiative efficiency of 4.3 %. This study highlights the indispensable contribution of robotic networks to early afterglow observations and advances our understanding of GRB 230204B unique characteristics and underlying jet physics.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are among the most energetic events in the Universe, emitting vast amounts of energy within a few seconds to minutes, often followed by an afterglow that can be observed across the electromagnetic spectrum (Piran, 1999; Kumar & Zhang, 2015). The discovery of afterglows in the late 1990s marked a significant milestone in GRB research, enabling precise localization and redshift determination, which, in turn, provided insights into the distances and energies involved (Costa et al., 1997; van Paradijs et al., 1997; Frail et al., 1997). The study of GRB afterglows, particularly in the optical regime, has been pivotal in understanding the progenitor systems, the environment surrounding the burst, and the mechanisms driving the emission (Kann et al., 2010; Li et al., 2012; Gupta, 2023). The temporal evolution of these afterglows, often characterized by a power-law decay, can reveal crucial details about the jet dynamics, energy injection, and interaction with the surrounding medium (Meszaros & Rees, 1997a; Sari et al., 1998).
@@
In recent times, robotic telescope networks, such as Mobile Astronomical System of Telescope-Robots (MASTER), Burst Observer and Optical Transient Exploring System (BOOTES), and others, have become indispensable in the rapid follow-up of GRBs, often identifying optical afterglows within minutes to hours (especially for Fermi Gamma-ray Burst Monitor (GBM) detected GRBs with large error circles) of the initial burst (Lipunov et al., 2010a; Castro-Tirado et al., 2012). Fermi GBM’s localization accuracy, while adequate for high-energy studies, often leaves large localization uncertainties spanning a few to tens of degrees radius, making rapid and wide-field follow-up observations essential. These robotic telescope networks, equipped with wide-field cameras and fully automated systems, are specifically designed to address this challenge. Their ability to quickly respond to GRB alerts and scan large areas of the sky allows them to pinpoint transient optical emissions within the Fermi error regions. Additionally, their prompt response times enable them to observe early afterglow phases, providing data at these early times critical for informing the onset and evolution of the emission (Gupta et al., 2021b; Ror et al., 2023b; Sánchez-Ramírez et al., 2024). While these observations are useful for constraining GRB redshifts indirectly by identifying afterglow candidates for follow-up spectroscopic studies, they typically rely on single-band detections and are limited in sensitivity, restricting their ability to detect only the bright afterglows. Despite these limitations, the robotic nature of these observatories ensures continuous and autonomous temporal coverage, improving the chances of detecting GRBs in real-time and providing crucial data to guide further multi-wavelength follow-up efforts (Castro-Tirado, 2023).
@@
GRB 230204B, detected by the Fermi GBM and MAXI, stands out as one of the most energetic GRBs observed, with an isotropic equivalent energy (Eγ,iso) exceeding 1054 erg. The optical afterglow of this burst was discovered by the MASTER and BOOTES robotic telescope networks approximately 1.3 ks post-burst, marking the beginning of the broadband follow-up campaign. Additionally, at the time of MASTER/BOOTES detection ( 1.3 ks post-burst), the optical luminosity was found to be exceptionally high when compared to other bright bursts, suggesting that GRB 230204B harbored an extremely luminous afterglow phase. In this paper, we present a detailed analysis of the optical and multi-wavelength follow-up observations (see Section 2) of GRB 230204B, including data from MASTER, BOOTES, Devasthal Optical Telescope (DOT), Swift Ultraviolet and Optical telescope (UVOT) and radio telescope. We explore the temporal and spectral characteristics of the prompt emission and afterglow and compare its optical luminosity with other known GRBs, the results are presented in Section 3. Discussion on potential physical mechanisms driving this energetic burst and the afterglow brightness comparison are given in Section 4. The key findings are summarized in Section 5.
@@ 2 Observations and data analysis
This section details the observation campaign for GRB 230204B, observed by multiple space-based and ground-based telescopes across a broad wavelength range, from gamma rays to radio frequencies. Comprehensive details of the observing campaign are provided below and the sequence of prompt and afterglow observations for GRB 230204B is shown in Figure 1.
@@ 2.1 Prompt Emission
GRB 230204B was first detected on 4th February 2023 by the Fermi satellite at 21:44:27.20 UT (hereafter T0), with duration of 216 seconds in 50-300 keV (Poolakkil et al., 2023). MAXI also detected it at 21:47:51 UT and the burst position was constrained to approximately RA = 13h 10m 19s, Dec = -21° 45′ 07′′ (J2000) with a statistical uncertainty region of 0.1°(Serino et al., 2023). The burst did not trigger the Neil Gehrels Swift Observatory; henceforth Swift (Gehrels et al., 2004). However, following the MAXI trigger notification, the Gamma-ray Urgent Archiver for Novel Opportunities (GUANO, Tohuvavohu et al. 2020) pipeline requested data from the onboard buffer of the Swift Burst Alert Telescope (BAT) and initiated a search within the 200-second event mode data centered around T0. The GRB was detected in this Swift/BAT GUANO data, determining its to be approximately 65 seconds (Kennea et al., 2023). The burst was also detected by the Indian multiwavelength satellite AstroSat, which measured a of seconds (Waratkar et al., 2023). The prompt emission of the burst was also seen by AGILE (Casentini et al., 2023) and GRBAlpha (Dafcikova et al., 2023) missions, with every mission providing a slightly different value of duration as expected due to the different sensitive and spectral coverage.
@@ 2.2 Prompt X-ray Emission: MAXI
MAXI (mounted on the ISS) scans a position of the sky every 92 minutes (Matsuoka et al., 2009). The position of the GRB moved into the MAXI field of view at T0 +192.8 sec, and the scan transit ended at T0 +248.8 sec. During the scan transit, the MAXI/GSC nova alert system (Negoro et al., 2016) triggered the burst at 21:47:51 UT (T0 + 203.8 sec), with MAXI’s detection revealing the soft nature of the burst. Assuming the source flux was constant over the transit, the X-ray flux in the 4.0–10.0 keV band was calculated to be 628 49 mCrab (1 sigma error, Serino et al. 2023).
After the detection of prompt emission, MAXI/GSC could not observe the position of the burst for about 6 hours due to the observation constraint of the radiation zone (Sugizaki et al., 2011) in the International Space Station (ISS) orbit. The observation resumed at T0 + 22.5 ks, and there were 11 consecutive scan transits before the start of the observations by Swift/XRT. The afterglow was not detected in any transit. We calculated 3 sigma upper limits in the 2–20 keV band (see Table 8 in appendix) for these transits following the method by Sugita et al. (2018).
@@ 2.3 Afterglow: Soft X-ray Emission
The fading X-ray counterpart of GRB 230204B was detected by Swift/XRT approximately 80 ks after the burst trigger (D’Elia et al., 2023). Given the faint nature of the X-ray afterglow at such a late epoch, all observations were performed in photon counting (PC) mode to optimize sensitivity for faint sources. Follow-up observations were conducted during two epochs (XRT exposure times of 5.1 ks, and 4.5 ks, respectively), on 5 February 2023 and 14 February 2023, resulting in only two usable data points (see Table A2 of the appendix). The final observation on February 14 approached the detection limit of the instrument, emphasizing the faintness of the X-ray emission and the challenges of constraining the X-ray afterglow’s properties at these late times.
@@ 2.4 Afterglow: Optical Emission
We discovered the optical afterglow of GRB 230204B using two robotic telescope networks (using MASTER in the unfiltered band and using BOOTES in the clear band, respectively) independently (see Figure 1). The optical counterpart of the burst was also identified by the 0.7m GROWTH-India Telescope (GIT) approximately 1.57 hours (later than MASTER and BOOTES) after the burst trigger (Swain et al., 2023). This optical afterglow was subsequently observed by several other ground-based telescopes (Smartt et al., 2023; Ror et al., 2023a). The redshift of the burst, = 2.142, was determined through spectroscopic observations using the X-shooter spectrograph on the Very Large Telescope (VLT, Saccardi et al. 2023). The given magnitude (for the positive detection) and upper limits within 3 are listed in Table LABEL:tab:optical.
@@ 2.4.1 MASTER robotic telescope
MASTER global network consists of 9 identical fully robotic wide-field twin telescopes with 4-8 square degrees field of view, distributed around the Earth for all-sky monitoring up to 19-20 mag with the goal of detecting optical transients, including high energy astrophysics events such as GRBs, Gravitational waves, ultra-high energy neutrino and Fast Radio Bursts (Lipunov et al., 2010b, 2019, 2022, 2023). We observed such events as soon as possible upon receiving the alerts (Ershova et al., 2020) using the GCN system (Barthelmy et al., 1998; Barthelmy, 2003, 2016) every night in both, the fully robotic open (8 square degrees FOV) and closed (4 square degrees FOV) modes, being ready for observations with fully clear hemisphere (clamshell dome constructions). Open mode is used in the case of alerts provided by Fermi, LIGO/Virgo, IceCube, ANTARES/KM3Net (with large error boxes) when at least 2 images per night are required if large error boxes are at least 5 degrees above the horizon, in order to cover as much as possible with our given field of view. Closed mode (using both tubes in parallel Kornilov et al. (2012)) is preferred for events with error boxes smaller than 1 degree such as the ones provided by Swift, MAXI, Fermi-LAT, GECAM, Einstein Probe, INTEGRAL, when polarization observations can be undertaken (Troja et al., 2017; Jordana-Mitjans et al., 2020; Gupta et al., 2022a). The first exposure duration also depends on the time between the trigger and the start of observation and can be increased from 10 sec at the first minute to 180 sec at later stages in order to detect the fainter objects.
GRB 230204B was observed by MASTER-Kislovodsk and MASTER-SAAO. First images were received for the MAXI error box, then telescopes automatically pointed to Fermi field. The MASTER-Kislovodsk robotic telescope started observing the MAXI trigger 979708795 observation at 2023-02-04 22:05:55 UT (1064 sec after trigger time which was 74 sec after the notification time, see Figure 1) at 11∘ altitude, in 44∘ distance to the Moon (0.3 phase, that creates significant interference for wide-field instruments observations) through clouds at horizon with =14.8 mag at first image in polarization filter.
@@
For the photometric analysis of this object, catalog data for comparison stars were sourced from Gaia EDR3 to evaluate the variance in magnitude errors. The photometry dataset excludes frames where the comparison star data were deemed unreliable. The magnitude error for comparison stars was computed as the deviation between the weighted average magnitude (where weights are the inverse square of the signal-to-noise ratio) and the magnitude measured on individual frames. Two approaches were employed to determine the stellar magnitude. In the first approach, a reference star from Gaia EDR3 was identified, its flux on the observational frame measured, and the magnitude calculated using Pogson’s scale. In the second approach, a zero-point correction was applied to the reference star’s magnitude, utilizing the comparison stars as secondary references. The second method generally produced a smaller magnitude variance compared to the first, owing to the improved calibration achieved through the zero-point correction. This dual-approach methodology ensures robust photometric accuracy while mitigating systematic errors.

### v2

# Extremely luminous optical afterglow of an energetic gamma-ray burst GRB 230204B

@@ Abstract
Robotic telescope networks play an important role in capturing early and bright optical afterglows, providing critical insights into the energetics and emission mechanisms of GRBs. In this study, we analyze GRB 230204B, an exceptionally energetic and multi-pulsed long GRB, detected by the Fermi GBM and MAXI detectors, with an isotropic equivalent gamma-ray energy exceeding 1054 erg. Time-resolved spectral analysis reveals a transition in the prompt emission from hard (sub-photospheric dominated) spectra during early pulses to softer (synchrotron radiation dominated) spectra in later pulses, indicative of a hybrid jet composition. We report the discovery and characterization of the optical afterglow using the MASTER and BOOTES robotic telescope networks, which enabled rapid follow-up observations starting at 1.3 ks post-burst. The optical luminosity at this time was exceptionally high, surpassing that of many other optically bright GRBs, such as GRB 990123, GRB 080319B, etc. This places the burst among the most luminous optical GRBs observed to date. Long-term radio observations extending to 335 days post-burst were conducted with the ATCA. Multi-wavelength modeling, incorporating data from MASTER, BOOTES, DOT, Swift/XRT, and radio observations, was conducted using an external ISM forward-shock top-hat jet model with afterglowpy. The results reveal a narrow and highly collimated jet with a circumburst density of 28.12 cm-3, kinetic energy 4.18 erg, and a relatively low value of = 2.14 , indicating shock-compression of magnetic field in the surrounding interstellar medium. We constrained a low radiative efficiency of 4.3 %. This study highlights the indispensable contribution of robotic networks to early afterglow observations and advances our understanding of GRB 230204B unique characteristics and underlying jet physics.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are among the most energetic events in the Universe, emitting vast amounts of energy within a few seconds to minutes, often followed by an afterglow that can be observed across the electromagnetic spectrum (1995ARA&A..33..415F; 1997ApJ...476..232M; 1998ApJ...497L..17S; 1999PhR...314..575P; 2000ApJ...543...66P; 2001ApJ...562L..55F; 2004RvMP...76.1143P; 2006RPPh...69.2259M; 2006ApJ...642..354Z; 2006ARA&A..44..507W; 2014ARA&A..52...43B; 2015PhR...561....1K; 2017ApJ...848L..12A). The discovery of afterglows in the late 1990s marked a significant milestone in GRB research. The first detection of an X-ray afterglow by 1997Natur.387..783C, optical afterglow by 1997Natur.386..686V, and radio afterglow by 1997Natur.389..261F enabled precise localization and redshift determination. This, in turn, provided insights into the distances and energies involved in GRBs (1999Natur.398..389K). The study of GRB afterglows, particularly in the optical regime, has been pivotal in understanding the progenitor systems (2001Sci...291...79M), the environment surrounding the burst (2010ApJ...720.1513K; 2012ApJ...758...27L), and the mechanisms driving the emission (2023arXiv231216265G). The temporal evolution of these afterglows, often characterized by a power-law decay, can reveal crucial details about the jet dynamics, energy injection, and interaction with the surrounding medium (1997ApJ...476..232M; 1998ApJ...497L..17S).
@@
In recent times, robotic telescope networks, such as Mobile Astronomical System of Telescope-Robots (MASTER, Li2022), Burst Observer and Optical Transient Exploring System (BOOTES, 2023NatAs...7.1136C), Las Cumbres Observatory (LCO) Global Telescope (2013PASP..125.1031B), Rapid Eye Mount (REM, 2004SPIE.5492.1613C), Liverpool Telescope (LT, 2016RMxAC..48...48C), Gravitational-wave Optical Transient Observer (GOTO, 2022MNRAS.511.2405S), Multicolor Imaging Telescopes for Survey and Monstrous Explosions (MITSuME, 2008AIPC.1000..543S), the Robotic Optical Transient Search Experiment (ROTSE III, 2003PASP..115..132A), Nazarbayev University Transient Telescope at Assy-Turgen Astrophysical Observatory (NUTTelA-TAO, 2022SPIE12184E..8AG), and others, have become indispensable in the rapid follow-up of GRBs, often identifying optical afterglows within minutes to hours (especially for Fermi Gamma-ray Burst Monitor (GBM) detected GRBs with large error circles) of the initial burst (2010AdAst2010E..30L; 2012ASInC...7..313C). Fermi GBM’s localization accuracy, while adequate for high-energy studies, often leaves large localization uncertainties spanning a few to tens of degrees radius (2009ApJ...702..791M; 2015ApJS..216...32C; 2020ApJ...895...40G), making rapid and wide-field follow-up observations essential. These robotic telescope networks, equipped with wide-field cameras and fully automated systems, are specifically designed to address this challenge. Their ability to quickly respond to GRB alerts and scan large areas of the sky allows them to pinpoint transient optical emissions within the Fermi error regions. Additionally, their prompt response times enable them to observe early afterglow phases, providing data at these early times critical for informing the onset and evolution of the emission (2021MNRAS.505.4086G; 2023ApJ...942...34R; 2024A&A...692A...3S). These observations are useful for constraining GRB redshifts indirectly by identifying afterglow candidates for follow-up spectroscopic studies. Furthermore, the robotic nature of these observatories ensures continuous and autonomous temporal coverage, improving the chances of detecting GRBs in real-time and providing crucial data to guide further multi-wavelength follow-up efforts (2023NatAs...7.1136C).
@@
GRB 230204B, detected by the Fermi GBM and MAXI, stands out as one of the most energetic GRBs observed, with an isotropic equivalent energy (Eγ,iso) exceeding 1054 erg (see Section 3.2). The rapid and precise localization of GRB 230204B by the MAXI detector triggered an immediate response from the MASTER and BOOTES robotic telescope networks, which provided crucial early-time optical data that significantly shaped our physical interpretation of this burst. These networks are specifically designed for the rapid follow-up of transient astronomical events, such as GRBs, and their automated observing capabilities are crucial for observing the early phases of the optical afterglow (2023NatAs...7.1136C). The MASTER network began observing GRB 230204B at 1.3 ks post-burst, providing the earliest optical detection of the afterglow of GRB 230204B and constrained the exceptionally high optical luminosity at the time of MASTER/BOOTES detection ( 1.3 ks post-burst), allowing us to place GRB 230204B among the most luminous optical afterglows known. The BOOTES network complemented these observations with additional early-time data, which were essential for constraining the decay phase of the optical light curve (see Section 3.3). In this paper, we present a detailed analysis of the optical and multi-wavelength follow-up observations (see Section 2) of GRB 230204B, including data from MASTER (2023GCN.33441....1L), BOOTES (2023NatAs...7.1136C), Devasthal Optical Telescope (DOT, 2023GCN.33284....1R; 2024BSRSL..93..683G), Swift Ultraviolet and Optical telescope (UVOT, 2023GCN.33292....1S; 2023Univ....9..113O), Growth India Telescope (GIT, 2023GCN.33269....1S), Asteroid Terrestrial-impact Last Alert System (ATLAS, 2023GCN.33278....1S), Very Large Telescope (VLT, 2023GCN.33281....1S), Australia Telescope Compact Array (ATCA, 2023GCN.33321....1G), and Australian Square Kilometre Array Pathfinder (ASKAP, 2020PASA...37...48M) telescope. We explore the temporal and spectral characteristics of the prompt emission and afterglow and compare its optical luminosity with other known GRBs, the results are presented in Section 3. Discussion on potential physical mechanisms driving this energetic burst and the afterglow brightness comparison are given in Section 4. The key findings are summarized in Section 5.
@@ 2 Observations and data analysis
This section details the observation campaign for GRB 230204B, observed by multiple space-based and ground-based telescopes across a broad wavelength range. Comprehensive details of the observing campaign are provided below, and the sequence of prompt and afterglow observations for GRB 230204B is shown in Figure 1.
@@ 2.1 Prompt Emission
GRB 230204B was first detected on 4th February 2023 by the Fermi GBM satellite at 21:44:27.20 UT (hereafter T0), with duration of 216 seconds in 50-300 keV (2023GCN.33288....1P). MAXI also detected it at 21:47:51 UT and the burst position was constrained to approximately RA = 13h 10m 19s, Dec = -21° 45′ 07′′ (J2000) with a statistical uncertainty region of 0.1°(2023GCN.33265....1S). The burst did not trigger the Neil Gehrels Swift Observatory; henceforth Swift (2004ApJ...611.1005G). However, following the MAXI trigger notification, the Gamma-ray Urgent Archiver for Novel Opportunities (GUANO, 2020ApJ...900...35T) pipeline requested data from the onboard buffer of the Swift Burst Alert Telescope (BAT) and initiated a search within the 200-second event mode data centered around T0. The GRB was detected in this Swift/BAT GUANO data, determining its to be approximately 65 seconds (2023GCN.33267....1K). The burst was also detected by the Indian multiwavelength satellite AstroSat, which measured a of seconds (2023GCN.33268....1W). The prompt emission of the burst was also seen by AGILE duration 145 sec, 2023GCN.33272....1C) and GRBAlpha ( of 207 sec, 2023GCN.33273....1D) missions, with every mission providing a slightly different value of duration as expected due to the different sensitive and spectral coverage.
@@ 2.2 Prompt X-ray Emission: MAXI
MAXI (mounted on the ISS) scans a position of the sky every 92 minutes (maxi). The position of the GRB moved into the MAXI field of view at T0 +192.8 sec, and the scan transit ended at T0 +248.8 sec. During the scan transit, the MAXI/GSC nova alert system (negoro) triggered the burst at 21:47:51 UT (T0 + 203.8 sec), with MAXI’s detection revealing the soft nature of the burst. The detection occurred during the final stages of the GRB emission, as evidenced by its temporal proximity to the main burst and the significant variability observed in the MAXI light curve during the scan transit (see Figure 4). The observed temporal variability is characteristic signature of prompt emission (2004IJMPA..19.2385Z; 2015PhR...561....1K; 2023A&A...671A.112C) rather than early afterglow, which typically exhibits smoother temporal evolution (1998ApJ...497L..17S; 2006ApJ...642..389N; 2006ApJ...647.1213O). These characteristics support the interpretation that the detected X-rays are part of the prompt emission of GRB 230204B rather than the afterglow. Assuming the source flux was constant over the transit, the X-ray flux in the 4.0–10.0 keV band was calculated to be 628 49 mCrab (1 sigma error, 2023GCN.33265....1S).
After the detection of the late phase of prompt emission of GRB 230204B by MAXI/GSC, it could not observe the position of the burst for about 6 hours due to the observation constraint of the radiation zone (sugizaki) in the International Space Station (ISS) orbit. The MAXI observations resumed at T0 + 22.5 ks, and there were 11 consecutive scan transits to search for an X-ray afterglow of GRB 230204B. The afterglow was not detected in any transit of MAXI observations. We calculated 3 sigma upper limits in the 2–20 keV band (see Table 9 in appendix) for these transits following the method by sugita. This non-detection by MAXI/GSC during the late phase further supports the prompt nature of the initial MAXI detection. A deeper search for the X-ray counterpart of GRB 230204B was later ( T0 + 80 ks) carried out by Swift/XRT (see Section 2.3).
@@ 2.3 Afterglow: Soft X-ray Emission
The fading X-ray counterpart of GRB 230204B was detected by Swift/XRT approximately 80 ks after the burst trigger (2023GCN.33285....1D). Given the faint nature of the X-ray afterglow at such a late epoch, all observations were performed in photon counting (PC) mode to optimize sensitivity for faint sources (2007A&A...469..379E; 2009MNRAS.397.1177E). Follow-up observations were conducted during two epochs (XRT exposure times of 5.1 ks, and 4.5 ks, respectively), on 5 February 2023 and 14 February 2023, resulting in only two usable data points (see Table A2 of the appendix). The final observation on February 14 approached the detection limit of the instrument, emphasizing the faintness of the X-ray emission and the challenges of constraining the X-ray afterglow’s properties at these late times.
@@ 2.4 Afterglow: Optical Emission
We discovered the optical afterglow of GRB 230204B using two robotic telescope networks (using MASTER in the unfiltered band and using BOOTES in the clear band, respectively) independently (see Figure 2). The optical counterpart of the burst was also identified by the 0.7m GROWTH-India Telescope (GIT) approximately 1.57 hours (later than MASTER and BOOTES) after the burst trigger (2023GCN.33269....1S). This optical afterglow was subsequently observed by several other ground-based telescopes (2023GCN.33278....1S; 2023GCN.33284....1R). Using the X-shooter spectrograph on the Very Large Telescope (VLT), 2023GCN.33281....1S estimated a redshift of = 2.142 for the burst. The given magnitude (for the positive detection) and upper limits within 3 are listed in Table LABEL:tab:optical.
@@ 2.4.1 MASTER robotic telescope
The MASTER global network comprises nine identical fully robotic wide-field twin telescopes, distributed around the Earth for all-sky monitoring up to 19-20 mag. Its primary goal is to detect optical transients, including high-energy astrophysical events such as GRBs (Li2010; Li2019; Li2022; Li2023). The network operates in both open (8 square degrees FOV) and closed (4 square degrees FOV) modes, enabling rapid follow-up observations of alerts from facilities such as Fermi, LIGO/Virgo, IceCube, ANTARES/KM3Net, Swift, MAXI, Fermi-LAT, GECAM, Einstein Probe, and INTEGRAL (Kornilov; Troja; Mundell; Gupta1).
GRB 230204B was observed by MASTER-Kislovodsk and MASTER-SAAO (South African Astronomical Observatory). First images were received for the MAXI error box, then telescopes automatically pointed to Fermi field. The MASTER-Kislovodsk robotic telescope started observing the MAXI trigger 979708795 observation at 2023-02-04 22:05:55 UT (1064 sec after trigger time which was 74 sec after the notification time, see Figure 2) at 11∘ altitude, in 44∘ distance to the Moon (0.3 phase, that creates significant interference for wide-field instruments observations) through clouds at horizon with =14.8 mag at first image in polarization filter. The transient was not detected in this initial image.
@@
For the photometric analysis of this object, catalog data for comparison stars were sourced from Gaia EDR3 to evaluate the variance in magnitude errors. The photometry dataset excludes frames where the comparison star data were deemed unreliable. The magnitude errors for comparison stars were computed as the deviation between the weighted average magnitude (where weights are the inverse square of the signal-to-noise ratio) and the magnitude measured on individual frames. Two approaches were employed to determine the stellar magnitude. In the first approach, a reference star from Gaia EDR3 was identified, its flux on the observational frame measured, and the magnitude calculated using Pogson’s scale. This method served as a validation step to assess consistency and identify potential systematic errors. In the second approach, which was used for the final magnitude determination, a zero-point correction was applied to the reference star’s magnitude, utilizing the comparison stars as secondary references. The second method generally produced a smaller magnitude variance compared to the first, owing to the improved calibration achieved through the zero-point correction. By employing this dual-approach methodology, we ensured robust photometric accuracy while mitigating systematic errors. The final results presented in this study are based solely on the second method.